# Connected Speech-Based Cognitive Assessment in Chinese and English

### v1

@@ Abstract
We present a novel benchmark dataset and prediction tasks for investigating approaches to assess cognitive function through analysis of connected speech. The dataset consists of speech samples and clinical information for speakers of Mandarin Chinese and English with different levels of cognitive impairment as well as individuals with normal cognition. These data have been carefully matched by age and sex by propensity score analysis to ensure balance and representativity in model training. The prediction tasks encompass mild cognitive impairment diagnosis and cognitive test score prediction. This framework was designed to encourage the development of approaches to speech-based cognitive assessment which generalise across languages. We illustrate it by presenting baseline prediction models that employ language-agnostic and comparable features for diagnosis and cognitive test score prediction. The models achieved unweighted average recall was 59.2% in diagnosis, and root mean squared error of 2.89 in score prediction.
@@ 1 Introduction
Cognitive problems such as memory loss, speech and language impairment, and reasoning difficulties occur frequently among older adults and often precede the onset of dementia syndromes. Due to the high prevalence of dementia and the costs this implies to health systems worldwide [1], research into cognitive impairment for the purposes of dementia prevention and early detection has become a priority in healthcare. There is a need for cost-effective and scalable methods for assessment of cognition and detection of impairment, from its most subtle forms to severe manifestations of dementia. Speech is an easily collectable behavioural signal which reflects cognitive function and, therefore, could potentially serve as a digital biomarker of cognitive function, presenting a unique opportunity for application of speech technology [2].
We aim to assess speech as a marker of cognition in a global health context by investigating its application to modelling cognitive health indicators in two major languages, namely, Chinese and English. In this paper, we focus on prediction of cognitive test scores and diagnosis of mild cognitive impairment (MCI) in older speakers of Chinese and English, using samples of connected speech. We are particularly interested in investigating approaches that are language independent or build on comparable features. To this end, we have created, and are sharing with the research community, a dataset comprising recorded speech from study participants carrying out picture description tasks along with clinical and neuropsychological test data.
This dataset can be used as a benchmark for speech processing and machine learning tasks that are relevant to the detection of cognitive decline through analysis of connected speech data. We hope that this new resource will stimulate research on speech biomarkers among members of the speech, signal processing, machine learning, natural language processing and biomedical research communities, enabling them to test existing methods or develop novel approaches on a new, standardised dataset which will remain available to the community for future research and replication of results.
@@ 2 Background
Clinical tests, such as the Mini-Mental State Examination (MMSE), are often part of these studies as mere data descriptors, rarely used in prediction. Some studies [8, 9, 10, 11] have used MMSE results as a baseline for classification, against which to compare the speech-based classifier, but very few available studies go beyond classification and use speech-based approaches to predict MMSE scores or other cognitive tests. However, there has been a shift of focus toward it in recent years. For instance, lexico-semantic features extracted from picture descriptions have been used in a model that was able to explain 51% of the variance of cognitive scores at the time of speech collection, and 56% of cognitive scores in a 12-month follow-up [12]. Other approaches have also been published, such as [13], which used BERT to predict MMSE scores from denoised speech recordings from picture description tasks and reported a root mean squared error (RMSE) of 3.76. Another study reported that acoustic features alone predict MMSE scores with a mean absolute error (MAE) of 5.66 and an $R^{2}$ of 0.125, with a linear regression analysis that improved by adding age, sex, and years of education to the model, yielding a MAE = 4.97 and $R^{2}$ = 0.261 [14] on the balanced dataset used by [7].
@@ 3 Data
The training set contained both Chinese and English samples with three picture descriptions per participant. The test set comprised recordings from different participants, with the same mix of languages and picture descriptions. Basic descriptive statistics of training and test set are shown in Table 1. Overall, there are 507 speech samples (261 Chinese and 246 English) with total duration of 528 minutes, ratio of training to test samples is just over 3:1. The dataset has been made available to the wider research community via DementiaBank [^1] https://dementia.talkbank.org/.
@@ 4 Cognitive assessment tasks
The benchmark presented in this paper encompasses the following tasks: (a) a classification task, where we aim to create models to distinguish NC speech from MCI speech, and (b) a cognitive test score prediction (regression) task, where we create models to infer the subject’s MMSE scores based on connected (spontaneous) speech data.
The MCI classification task will be evaluated through specificity ($\sigma$), sensitivity ($\rho$) and $F_{1}$ scores for the MCI category. These metrics will be computed as follows: $\sigma=\frac{T_{N}}{T_{N}+F_{P}}$, $F_{1}=\frac{2\pi\rho}{\pi+\rho}$, where $\pi=\frac{T_{P}}{T_{P}+F_{P}}$, $\rho=\frac{T_{P}}{T_{P}+F_{N}}$, $N$ is the number of patients, $T_{P}$ is the number of true positives, $T_{N}$ is the number of true negatives, $F_{P}$ is the number of false positives and $F_{N}$ the number of false negatives. The balanced accuracy metric (unweighted average recall, UAR) will be used for the overall ranking of this task’s results. It is defined as follows: $\operatorname{UAR}=\frac{\sigma+\rho}{2}$.
The MMSE regression task will be assessed using the RMSE, defined as $\operatorname{RMSE}=\sqrt{\frac{\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}}{N}}$, where $\hat{y}$ is the predicted MMSE score, $y$ is the patient’s actual MMSE score, and $\bar{y}$ is the mean score.
@@ 5.1 Feature extraction
The feature extraction procedure aimed to identify speech features that could generalise well across the two languages. For acoustic features, we tested two different approaches: a traditional feature engineering approach with a feature set that has been found useful in emotion recognition and other computational paralinguistics tasks (eGeMAPs), and a self-supervised feature learning approach (wav2vec).
@@
For self-supervised feature extraction we used the pre-trained model wav2vec and extracted features directly from raw audio [30]. To balance the duration of all audio recordings, we zero-padded the audio recordings for feature extraction. Next, we applied a dropout layer, followed by a feature aggregation layer and another dropout layer. For dimensionality reduction, we used MaxPool1d layer (with a size of 42000, and a stride of 10,000). The result was used as input features for the multilayer perceptron (MLP) models. This resulted in 512 features per audio recording.
@@ 6 Results
For the classification (diagnostic) task, our model achieved a test-data UAR of 59.18% while fusing the wav2vec and eGeMAPs features. The full set of results is shown in Table 2. Confidence interval were obtained through bootstrapping over 1000 runs [31]. The baseline result for this task is 59.18% UAR obtained on test data ($\sigma=0.587$, $\rho=0.597$, $\pi=0.617$). The overall accuracy score was 0.592, while F1 reached 0.602. The confusion matrix for test-set classification with the feature fusion model is shown in Figure 2. Figure 3 shows the effect of each feature set on classification performance. color=yellow]SL: The problem with these results is that if we were to choose a model purely based on CV we would have chosen the wav2vec, which clearly overfits, giving a test set UAR of just over 46%. We need to address this.
Figure 2: Confusion matrix for fusion test results.
@@
For the regression task, the comparable linguistic features on their own proved to be the most effective features, with RMSE scores of 2.86 and 2.89 for validation and test sets, respectively. Combining wav2vec and linguistic features also proved effective, but the eGeMAPs acoustic features were not found to be useful in this task.
@@ 7 Discussion
The present dataset is considerably less heterogeneous in terms of diagnoses and cognitive test scores than most public data used to date in research on predictive models for cognitive function assessment, including the few existing cross- and multi-lingual speech datasets used in this area [18, 16]. This makes the learning tasks defined in this paper harder, as they need to discriminate over a narrower range of values. However, the performance of our baseline models is comparable to that of those models.
For the cognitive score prediction task (regression), we achieve an RMSE score of 2.89, while [13] reports an RMSE of 3.76, but using only the English subset of our data. The most comparable research is that conducted in a signal processing grand challenge to generalise speech-based predictive models across Greek and English [18, 32]. The best performing models achieved a classification accuracy between 69% and 87% (AD vs NC) and a RMSE between 4.79 and 3.72 for MMSE score prediction. However, these models involved training in one language and testing in another, while our experimental setup yields comparable results combining both languages for training and tests sets, to our knowledge, for the first time.
Our goal was not to push the state-of-the-art on these datasets, but rather to establish proof-of-concept for our multilingual speech-only model’s capabilities to predict MMSE scores and detect MCI on a homogeneous multilingual dataset. Given that (a) our baseline results are comparable to other models in the literature (59.18%UAR and 2.89 RMSE), (b) both MMSE prediction and MCI detection are significantly sparse compared to AD detection in the literature [2], and (c) this is the first speech-only model built through combining datasets in two different languages, we argue that our work significantly contributes to the development of the field and will serve as a workable baseline for the wider research community.
As a limitation of this study that is generalised to the research field, it should be noted that MMSE has been criticised for low discrimination, especially in preclinical dementia [33]. Therefore, future studies should aim to focus on other cognitive tests, more able to discriminate early stages of cognitive impairment.
@@ 9 Acknowledgements
The authors would like to thank ISCA and the organising committees of past Interspeech conferences for their help and for kindly providing the previous version of this template.

### v2

@@ Abstract
We present a novel benchmark dataset and prediction tasks for investigating approaches to assess cognitive function through analysis of connected speech. The dataset consists of speech samples and clinical information for speakers of Mandarin Chinese and English with different levels of cognitive impairment as well as individuals with normal cognition. These data have been carefully matched by age and sex by propensity score analysis to ensure balance and representativity in model training. The prediction tasks encompass mild cognitive impairment diagnosis and cognitive test score prediction. This framework was designed to encourage the development of approaches to speech-based cognitive assessment which generalise across languages. We illustrate it by presenting baseline prediction models that employ language-agnostic and comparable features for diagnosis and cognitive test score prediction. Unweighted average recall was 59.2% in diagnosis, and root mean squared error was 2.89 in score prediction.
@@ 1 Introduction
Cognitive problems such as memory loss, speech and language impairment, and reasoning difficulties occur frequently among older adults and often precede the onset of dementia syndromes. Due to the high prevalence of dementia and the costs this implies to health systems worldwide [1], research into cognitive impairment for the purposes of dementia prevention and early detection has become a priority in healthcare. There is a need for cost-effective and scalable methods for assessment of cognition and detection of impairment, from its most mild forms to severe manifestations of dementia. Speech is an easily collectable behavioural signal which reflects cognitive function and, therefore, could potentially serve as a digital biomarker of cognitive function, presenting a unique opportunity for application of speech technology [2].
We aim to assess speech as a behavioural marker of cognition in a global health context by investigating its application to the modelling of cognitive health indicators in two major languages, namely, Chinese and English. In this paper, we focus on prediction of cognitive test scores and diagnosis of mild cognitive impairment (MCI) in older speakers of Chinese and English, using samples of connected speech produced in picture description tasks. Our aim was to investigate approaches that are language independent or build on comparable features. To this end, we created, and are sharing with the research community, recorded speech from study participants doing picture description tasks along with clinical and neuropsychological test data.
This dataset has been used as a benchmark for speech processing and machine learning tasks that are relevant to the detection of cognitive decline through analysis of connected speech data. It formed the basis of the TAUKADIAL Challenge, at Interspeech 2024 (http://luzs.gitlab.io/taukadial/). We hope that this new resource will stimulate research on speech biomarkers in the speech, signal processing, machine learning and biomedical research communities, enabling them to test existing methods or develop novel approaches on a new, standardised dataset which will remain available to the community for future research and replication of results.
@@ 2 Background
Clinical tests such as the Mini-Mental State Examination (MMSE) are often part of these studies as mere data descriptors, rarely used in prediction. Some studies [8, 9, 10, 11] have used MMSE results as a baseline for classification, against which to compare the speech-based classifier, but very few available studies go beyond classification and use speech-based approaches to predict MMSE scores or other cognitive tests. However, there has been a shift of focus toward it in recent years. For instance, lexico-semantic features extracted from picture descriptions have been used in a model that was able to explain 51% of the variance of cognitive scores at the time of speech collection, and 56% of cognitive scores in a 12-month follow-up [12]. Other approaches have also been published, such as [13], which used BERT to predict MMSE scores from denoised speech recordings from picture description tasks and reported a root mean squared error (RMSE) of 3.76. Another study reported that acoustic features alone predict MMSE scores with a mean absolute error (MAE) of 5.66 and an $R^{2}$ of 0.125, with a linear regression analysis that improved by adding age, sex, and years of education to the model, yielding a MAE = 4.97 and $R^{2}$ = 0.261 [14] on the balanced dataset used by [7].
@@ 3 Data
The training set contained both Chinese and English samples with three picture descriptions per participant. The test set comprised recordings from different participants, with the same mix of languages and picture descriptions. Basic descriptive statistics of training and test set are shown in Table 1. Overall, there are 507 speech samples (261 Chinese and 246 English) with total duration of 528 minutes, ratio of training to test samples is just over 3:1. The dataset has been made available to the wider research community via DementiaBank [21].
@@ 4 Cognitive assessment tasks
The benchmark presented in this paper encompasses the following tasks: (a) a classification task, where we aimed to create models to distinguish NC speech from MCI speech, and (b) a cognitive test score prediction (regression) task, where we created models to infer the subject’s MMSE scores based on connected (spontaneous) speech data.
The MCI classification task is evaluated through specificity ($\sigma$), sensitivity ($\rho$) and $F_{1}$ scores for the MCI category. These metrics are computed as follows: $\sigma=\frac{T_{N}}{T_{N}+F_{P}}$, $F_{1}=\frac{2\pi\rho}{\pi+\rho}$, where $\pi=\frac{T_{P}}{T_{P}+F_{P}}$, $\rho=\frac{T_{P}}{T_{P}+F_{N}}$, $N$ is the number of patients, $T_{P}$ is the number of true positives, $T_{N}$ is the number of true negatives, $F_{P}$ is the number of false positives and $F_{N}$ the number of false negatives. The balanced accuracy metric (unweighted average recall, UAR) is used for the overall ranking of this task’s results. It is defined as follows: $\operatorname{UAR}=\frac{\sigma+\rho}{2}$.
The MMSE regression task is assessed using the RMSE, defined as $\operatorname{RMSE}=\sqrt{\frac{\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}}{N}}$, where $\hat{y}$ is the predicted MMSE score, $y$ is the patient’s actual MMSE score, and $\bar{y}$ is the mean score.
@@ 5.1 Feature extraction
The feature extraction procedure aimed to identify speech features that could generalise well across the two languages. For acoustic features, we tested two different approaches: a traditional feature engineering approach with a feature set that has been found useful in emotion recognition and other computational paralinguistics tasks (eGeMAPs), and a self-supervised feature learning approach.
@@
For self-supervised feature extraction we used the pre-trained model wav2vec, without fine tuning, and extracted features directly from raw audio [30].
To balance the duration of all audio recordings, we zero-padded the audio recordings for feature extraction. The features are extracted from the feature extractor layer. Then we applied a dropout layer, followed by a feature aggregation layer and another dropout layer. For dimensionality reduction, we used MaxPool1d layer (with a size of 42000, and a stride of 10,000). The result was used as input features for the multilayer perceptron (MLP) models. This resulted in 512 features per audio recording.
@@ 6 Results
For the classification (diagnostic) task, our model achieved a test-data UAR of 59.18% while fusing the wav2vec and eGeMAPs features. The full set of results is shown in Table 2. Confidence interval were obtained through bootstrapping over 1000 runs [31]. The baseline result for this task is 59.18% UAR obtained on test data ($\sigma=0.587$, $\rho=0.597$, $\pi=0.617$). The results were very similar in both languages (English: UAR$=60.00\%$, $\sigma=0.40$, $\rho=0.80$; Chinese: UAR$=60.04\%$, $\sigma=0.39$, $\rho=0.81$). The overall accuracy score was 0.592, while F1 reached 0.602. Figure 2 shows the effect of each feature set on classification performance and Table 3 shows the results for both languages for the best performing methods. color=yellow]SL: The problem with these results is that if we were to choose a model purely based on CV we would have chosen the wav2vec, which clearly overfits, giving a test set UAR of just over 46%. We need to address this.
@@
For the regression task, the comparable linguistic features on their own proved to be the most effective features, with RMSE scores of 2.86 (r = 0.514) and 2.89 (r = 0.337) for validation and test sets, respectively. Combining wav2vec and linguistic features also proved effective, but the eGeMAPs acoustic features were not found to be useful in this task. Unlike classification, regression results differed by language. For English, the RMSE was 1.75, while for Chinese the RMSE was 3.71, reflecting the standard deviations of MMSE (4.11 for Chinese and 1.27 for English).
@@ 7 Discussion
The present dataset is considerably less heterogeneous in terms of diagnoses and cognitive test scores than most public data used to date in research on predictive models for cognitive function assessment, including the few existing cross- and multi-lingual speech datasets used in this area [18, 16]. This makes the learning tasks defined in this paper harder, as they need to discriminate over a narrower range of values. However, our baseline models perform comparably to those models.
For the cognitive score prediction task (regression), we achieve an RMSE score of 2.89, while [13], for instance, reports an RMSE of 3.76, but using only the English subset of our data. The most comparable research is that conducted in a signal processing grand challenge to generalise speech-based predictive models across Greek and English [18, 32]. The best performing models achieved a classification accuracy between 69% and 87% (AD vs NC) and a RMSE between 4.79 and 3.72 for MMSE score prediction. However, these models involved training in one language and testing in another, while our experimental setup yields comparable results combining both languages in training and test, to our knowledge, for the first time.
Our goal was not to push the state-of-the-art on these datasets, but rather to establish proof-of-concept for our multilingual speech-only model’s capabilities to predict MMSE scores and detect MCI on a homogeneous multilingual dataset. Given that (a) our baseline results are comparable to other models in the literature (59.18%UAR and 2.89 RMSE), (b) both MMSE prediction and MCI detection are relatively uncommon compared to AD detection in the literature [2], (c) this is the first speech-only model built through combining datasets in two different languages, and (d) the models seem to generalise well for diagnosis across language (as suggested by the similarity in performance across the classification tasks), we argue that our work significantly contributes to the development of the field and will serve as a workable baseline for the wider research community.
As a limitation of this study, it should be noted that MMSE has been criticised for low discrimination (ceiling effect), especially in preclinical dementia [33]. This limitation is also common in similar studies in this research field. Therefore, future studies should aim to focus on other cognitive tests, more able to discriminate early stages of cognitive impairment.
@@
Table 3: Results Insights: Comparison of the best performing methods for the classification task (T.1) in % UAR and the MMSE regression task (T.2) in RMSE across both languages
T.1 T.2
English Cross-validation 52.7 1.43
Test 60.0 1.75
Chinese Cross-validation 48.3 3.73
Test 60.4 3.71
@@ 9 Acknowledgements
SL acknowledges support from the British Academy/Leverhulme through Research Grant number SRG2223/231718.